Question #1 The following selected data are taken from the financial statements of Chappelle Inc. The data are in alphabetical order 52,000 65.00 Shareholders equity. 310,900 4,000 Instructions Calculate the following ratios: (Hint: Remember: Total Assets Total Liabilities Shareholders equity) (a) Current ratio (b) Profit Margin (c) Basic earnings per share (d) Debt to total assets
Answer #1

a) Current Ratio = Current Assets/Current Liabilities

= Current Assets= Accounts Receivable+ Cash

=84,500+147,200= 231,700

Current Liabilities = Accounts Payable+ Salaries Payable+ Other Current Liabilities

=52,000+19,600+22,500

=94,100

Current Ratio = 231,700/94,100

=2.46

b) Profit Margin= Net Income/Sales

=174,000/590,000

=29.49%

c)Basic Earnings Per Share = Net income/Weighted Average Number of Common Shares

=174,000/4000

=43.50

d)Debt to total Assets = Debt/Total Assets

Debt= Total Assets- Current Liabilites- Shareholders Equity

=500,000-94,100-310,900

=95,000

Debt to total Assets =95,000/500,000

=19%